# Relevant or Random: Can LLMs Truly Perform Analogical Reasoning?

### v2

@@ Abstract
Analogical reasoning is a unique ability of humans to address unfamiliar challenges by transferring strategies from relevant past experiences. One key finding in psychology is that compared with irrelevant past experiences, recalling relevant ones can help humans better handle new tasks Gentner and Smith (2012). Coincidentally, the NLP community has also recently found that self-generating relevant examples in the context can help large language models (LLMs) better solve a given problem than hand-crafted prompts Yasunaga et al. (2024). However, it is yet not clear whether relevance is the key factor eliciting such capability, i.e., can LLMs benefit more from self-generated relevant examples than irrelevant ones? In this work, we systematically explore whether LLMs can truly perform analogical reasoning on a diverse set of reasoning tasks. With extensive experiments and analysis, we show that self-generated random examples can surprisingly achieve comparable or even better performance, e.g., 4 performance boost on GSM8K with random biological examples. We find that the accuracy of self-generated examples is the key factor and subsequently design two improved methods with significantly reduced inference costs. Overall, we aim to advance a deeper understanding of LLM analogical reasoning and hope this work stimulates further research in the design of self-generated contexts.
@@
Chengwei Qin††thanks: Equal contribution, order decided by coin flip.†, Wenhan Xia11footnotemark: 1♣, Tan Wang11footnotemark: 1†, Fangkai Jiao†, Yuchen Hu†, Bosheng Ding†, Ruirui Chen, Shafiq Joty†♠ †Nanyang Technological University ♣Princeton University ♠Salesforce Research Institute of High Performance Computing (IHPC), Agency for Science, Technology and Research (A*STAR) 1 Fusionopolis Way, #16-16 Connexis, Singapore 138632, Republic of Singapore
@@ 1 Introduction
With the recent advancements in scaling up model size and data, LLMs have demonstrated impressive zero-shot and few-shot performance across various reasoning tasks, especially, through advanced prompting methods like chain-of-thought (CoT) (Wei et al., 2022). Compared to common approaches such as zero or few-shot CoT (Zhou et al., 2022; Kojima et al., 2022; Wang et al., 2022b), Yasunaga et al. (2024) introduce LLM analogical reasoning, i.e., LLMs self-generate examples relevant to the query as context to better solve new problems; see Fig. 1 for an example. However, it remains unclear whether relevance is the key to eliciting such capability in LLMs. While several studies explore the influence of the relevance of demonstrations in in-context learning and CoT (Liu et al., 2022; Kim et al., 2022; Lyu et al., 2023; Chen et al., 2023; Yang et al., 2023; Wang et al., 2023; Yasunaga et al., 2024), none of them investigate whether self-generated relevant examples consistently outperform irrelevant ones in LLM analogical reasoning.
In this paper, to systematically assess the capability of LLMs to perform analogical reasoning, we conduct a series of ablation experiments on a variety of reasoning tasks including problems from GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and BIG-Bench Hard (BBH) (Suzgun et al., 2022). With extensive experiments, we aim to address the following two research questions:
@@
To answer these questions, we empirically analyze the analogical reasoning abilities of GPT-3.5 (turbo) and Llama series models (Touvron et al., 2023). Surprisingly, experimental results show that prompting LLMs to self-generate random examples can achieve comparable or even better performance on certain tasks which is not in line with the key claim of analogical reasoning in Gentner and Smith (2012), indicating that LLMs cannot always perform analogical reasoning. As for Q2, we point out through controlled experiments that the key factor is the accuracy of self-generated examples. Informed by these findings, we design two approaches that can outperform existing methods with significantly reduced inference costs. Specifically, we ask LLMs to randomly generate a few problems and manually verify their correctness, then use this fixed set of problems as in-context learning demonstrations for all test samples. Consistent observations across different model types consolidate the conclusions. We summarize the major contributions of our work below:
@@ 2.1 Chain-of-Thought Prompting
Chain-of-thought (CoT) prompting induces LLMs to generate intermediate reasoning steps before generating the final answer (Wei et al., 2022), greatly improving the reasoning capabilities of LLMs. Typical CoT prompting approaches include few-shot CoT (Wei et al., 2022; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a), taking several labeled demonstrations of the reasoning process, and zero-shot CoT, comprising only instructions like “Let’s think step by step” (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023). Other ongoing research on CoT has also explored (i) optimizing the demonstration selection (Fu et al., 2022; Lu et al., 2022; Qin et al., 2023), (ii) optimizing the quality of reasoning chains (Khot et al., 2022; Chen et al., 2022; Zhao et al., 2023; Shinn et al., 2023), and (iii) CoT in smaller language models (Magister et al., 2022; Ho et al., 2022; Fu et al., 2023).
@@ 2.2 LLM Analogical Reasoning
While few-shot CoT can provide more detailed reasoning guidance, it requires labeled examples which can be unavailable for a new task. To tackle this problem, Yasunaga et al. (2024) propose analogical prompting to guide LLMs to self-generate relevant exemplars as few-shot demonstrations, which is similar to analogical reasoning, i.e., humans can address new problems by drawing analogy from relevant past experience (Vosniadou and Ortony, 1989; Holyoak, 2012). In this work, we step forward to explore the intrinsic principle of LLM analogical reasoning. Specifically, we aim to investigate whether LLMs can authentically exhibit such reasoning capabilities and determine the extent to which the relevance of self-generated examples contributes to enhancing this process.
@@ 3 Methodology
We mainly follow the analogical prompting approach outlined in Yasunaga et al. (2024). Specifically, for a given target problem , analogical prompting introduces instructions like:
@@
In this work, we make minimal changes to the original prompt to guide LLMs to generate different types of irrelevant examples as context; see Fig. 2 for example prompts:
@@ 4.1 Experimental Setup
We mainly use GPT-3.5 (gpt-3.5-turbo) as the LLM and obtain all outputs from it with the temperature set to 0. We ask the LLM to self-generate 5 examples for GSM8K, 3 examples for MATH and BBH following Yasunaga et al. (2024).
@@ 4.2 Main Results
Problems in MATH span various subjects and difficulty levels. To investigate whether the inferior performance of relevant examples on MATH is accidentally caused by certain categories, we further report the accuracy across different subjects and difficulty levels in Table 3 and Fig. 3. The consistent performance gap between ‘relevant’ and other methods across different problem categories demonstrates the inherent flaws of relevant examples, indicating that mathematical reasoning tasks exhibit different analogical reasoning paradigms from other reasoning tasks.
@@
To quantitatively measure the relevance between the generated examples and the query, we compute the average cosine similarity between them. Following Zhang et al. (2023), we use the sentence transformer (Reimers and Gurevych, 2019) to encode all samples. For each method, the reported result is averaged across three seeds.
@@ 4.3 Further Analysis
Repeating Problems
Robustness to Prompt Formats
To verify the robustness of different methods to prompt formats, we experiment with a new prompt paraphrased from the original one by GPT-4 and present the results on GSM8K in Table 10. We also observe better performance with irrelevant examples than relevant ones, showing the robustness.
@@
Our experiments and analysis so far used GPT-3.5 as the LLM, which is closed-source and gets updated over time. To verify whether the observations and conclusions are consistent across different models and additionally for reproducibility, we extend the experiments to Llama-2-Chat (Touvron et al., 2023). Specifically, we use vLLM (Kwon et al., 2023) to serve a Llama-2-70b-Chat model for experiments and report the results of different methods/variants on GSM8K in Table 11 and Table 12. We can draw similar observations: (i) self-generated relevant examples underperform all types of irrelevant ones, (ii) ‘GPT4-Calibration’ consistently outperforms the other two variants, and (iii) and perform better than other approaches, demonstrating that the conclusions can be generalized to different models.
We further conduct experiments with the latest Llama-3-8B-Instruct model. The results reported in Table 11 demonstrate the generalizability of the conclusions across different model scales. In addition, since investigating analogical reasoning requires LLMs to self-generate different types of problems, we only experiment with instruction-tuned LLMs to ensure that they can follow the given instructions.
Different Numbers of Demonstrations
| Number | Relevant | N/A | |||
| 3 | 73.1 | 77.3 | 75.0 | 75.3 | 75.5 |
| 5 | 71.5 | 75.5 | 75.1 | 76.3 | 75.3 |
@@ 5 Conclusion
This work has several limitations. First, due to the inference cost of ChatGPT111experiments done between 12/2023 and 05/2024, we conduct experiments on subsets of the test data for mathematical reasoning tasks. Besides, we include 3 datasets requiring different reasoning capabilities in this work. A further improvement could be to explore more diverse types of tasks.
@@ A.2 Detailed Results for Different Random Seeds
A.3 Prompts and Outputs for Example Generation
A.4 Guided Problem Generation

### v3

@@ Abstract
Analogical reasoning is a unique ability of humans to address unfamiliar challenges by transferring strategies from relevant past experiences. One key finding in psychology is that compared with irrelevant past experiences, recalling relevant ones can help humans better handle new tasks. Coincidentally, the NLP community has also recently found that self-generating relevant examples in the context can help large language models (LLMs) better solve a given problem than hand-crafted prompts. However, it is yet not clear whether relevance is the key factor eliciting such capability, i.e., can LLMs benefit more from self-generated relevant examples than irrelevant ones? In this work, we systematically explore whether LLMs can truly perform analogical reasoning on a diverse set of reasoning tasks. With extensive experiments and analysis, we show that self-generated random examples can surprisingly achieve comparable or even better performance on certain tasks, e.g., 4 performance boost on GSM8K with random biological examples. We find that the accuracy of self-generated examples is the key factor and subsequently design two novel methods with improved performance and significantly reduced inference costs. Overall, we aim to advance a deeper understanding of LLM analogical reasoning and hope this work stimulates further research in the design of self-generated contexts.
@@
Chengwei Qin✦†††thanks: Equal contribution, order decided by coin flip., Wenhan Xia♣11footnotemark: 1, Tan Wang†11footnotemark: 1, Fangkai Jiao†, Yuchen Hu†, Bosheng Ding†, Ruirui Chen, Shafiq Joty†♠ ✦The Hong Kong University of Science and Technology (Guangzhou) ♣Princeton University †Nanyang Technological University ♠Salesforce Research Institute of High Performance Computing (IHPC), Agency for Science, Technology and Research (A*STAR), Singapore
@@ 1 Introduction
With the recent advancements in scaling up model size and data, LLMs have demonstrated impressive zero-shot and few-shot performance across various reasoning tasks, especially, through advanced prompting methods like chain-of-thought (CoT) (Wei et al., 2022). Compared to common approaches such as zero or few-shot CoT (Zhou et al., 2022; Kojima et al., 2022; Zhang et al., 2023a), Yasunaga et al. (2024) introduce LLM analogical reasoning, i.e., LLMs self-generate examples relevant to the query as context to better solve new problems; see Fig. 1 for an example. However, it remains unclear whether relevance is the key to eliciting such capability in LLMs. While several studies explore the influence of the relevance of demonstrations in in-context learning (ICL) and CoT (Liu et al., 2022; Kim et al., 2022; Lyu et al., 2023; Chen et al., 2023; Yang et al., 2023; Wang et al., 2023a; Alkhamissi et al., 2023; Yasunaga et al., 2024; Luo et al., 2024), none of them investigate whether self-generated relevant examples consistently outperform irrelevant ones in LLM analogical reasoning.
In this paper, to systematically assess the capability of LLMs to perform analogical reasoning, we conduct a series of ablation experiments on a variety of reasoning tasks including problems from GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and BIG-Bench Hard (BBH) (Suzgun et al., 2022). Furthermore, we evaluate the generalizability of our findings to other reasoning tasks, e.g., GPQA (Rein et al., 2024), in Section 4.3. With extensive experiments, we aim to address the following two research questions:
@@
To answer these questions, we empirically analyze the analogical reasoning abilities of GPT-3.5 (turbo), GPT-4o-mini, the Llama series (Touvron et al., 2023), and Qwen 2.5 (Yang et al., 2024) models. Surprisingly, experimental results show that prompting LLMs to self-generate random examples can achieve comparable or even better performance on certain tasks which is not in line with the key claim of analogical reasoning in Gentner and Smith (2012), indicating that LLMs cannot always perform analogical reasoning. As for Q2, we point out through controlled experiments that the key factor is the accuracy of self-generated examples. Informed by these findings, we design two approaches that can outperform existing methods with significantly reduced inference costs. Specifically, we ask LLMs to randomly generate a few problems and manually verify their correctness, then use this fixed set of problems as in-context learning demonstrations for all test samples. Consistent observations across different model types and scales consolidate the conclusions. We summarize the major contributions of our work below:
@@
Building on the findings, we propose two novel ICL-based approaches that improve performance while significantly reducing inference costs.
@@ 2.1 Chain-of-Thought Prompting
Chain-of-thought (CoT) prompting induces LLMs to generate intermediate reasoning steps before generating the final answer (Wei et al., 2022), greatly improving the reasoning capabilities of LLMs. Typical CoT prompting approaches include few-shot CoT (Wei et al., 2022; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a), taking several labeled demonstrations of the reasoning process, and zero-shot CoT, comprising only instructions like “Let’s think step by step” (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023a). Other ongoing research on CoT has also explored (i) optimizing the demonstration selection (Fu et al., 2022; Li and Qiu, 2023; Qin et al., 2024), (ii) optimizing the quality of reasoning chains (Khot et al., 2022; Chen et al., 2022; Shinn et al., 2023; Zhao et al., 2023; Besta et al., 2024), and (iii) CoT in smaller models (Magister et al., 2022; Ho et al., 2022; Fu et al., 2023; Qin et al., 2023; Ranaldi and Freitas, 2024; Peng et al., 2025).
@@ 2.2 LLM Analogical Reasoning
While few-shot CoT can provide more detailed reasoning guidance, it requires labeled examples which can be unavailable for a new task. To tackle this problem, Yasunaga et al. (2024) propose analogical prompting to guide LLMs to self-generate relevant exemplars as few-shot demonstrations, which is similar to analogical reasoning, i.e., humans can address new problems by drawing analogy from relevant past experience (Vosniadou and Ortony, 1989; Holyoak, 2012). LBS3 (Luo et al., 2024) explores curriculum learning which can better reflect human learning habits. In this work, we step forward to explore the intrinsic principle of LLM analogical reasoning. Specifically, we aim to investigate whether LLMs can authentically exhibit such reasoning capabilities and determine the extent to which the relevance of self-generated examples contributes to enhancing this process.
@@ 3 Methodology
Our analysis is based on the analogical prompting approach outlined in Yasunaga et al. (2024). Specifically, for a given target problem , analogical prompting introduces instructions like:
@@
In this work, we leverage similar prompts111Since our work aims to comprehensively explore and analyze the intrinsic principle of LLM analogical reasoning proposed in Yasunaga et al. (2024), we should follow the original design of the instructions to have a fair comparison and reliable analysis. to guide LLMs to generate different types of irrelevant examples as context; see Fig. 2 for example prompts:
@@ 4.1 Experimental Setup
We mainly use GPT-3.5 (gpt-3.5-turbo) as the LLM (see Section A.3 for more results with GPT-4o-mini) and obtain all outputs from it with the temperature set to 0. We ask the LLM to self-generate 5 examples for GSM8K, 3 examples for MATH and BBH following Yasunaga et al. (2024).
@@ 4.2 Main Results
Problems in MATH span various subjects and difficulty levels. To investigate whether the inferior performance of relevant examples on MATH is accidentally caused by certain categories, we further report the accuracy across different subjects and difficulty levels in Table 3 and Fig. 3. The consistent performance gap between ‘relevant’ and other methods across different problem categories demonstrates the inherent flaws of relevant examples, indicating that mathematical reasoning tasks exhibit different analogical reasoning paradigms from BBH.
@@
To quantitatively measure the relevance between the generated examples and the query, we compute the average cosine similarity between them. Following Zhang et al. (2023a), we use the sentence transformer (Reimers and Gurevych, 2019) to encode all samples. For each method, the reported result is averaged across three seeds (see Section A.4 for the decomposition of relevance).
@@ 4.3 Further Analysis
Difference from Previous Work
Apart from the comprehensive analysis, we have designed two novel ICL-based approaches that are completely different from the one in Yasunaga et al. (2024) (Q2-1). The difference lies mainly in the following two aspects: (i) The key claim in Yasunaga et al. (2024) is that we should guide the model to self-generate relevant examples as context. Motivated by the analysis and findings in our work (Q1 and Q2), our methods focus on ensuring the accuracy of self-generated examples rather than their relevance, which leads to better performance regardless of the problem type. (ii) As we have demonstrated that the relevance of self-generated examples does not matter, there is no need to generate relevant examples for each test query (the original method in Yasunaga et al. (2024)). In contrast, our methods use a fixed set of examples for all test queries, which significantly reduces the inference cost.
| Dataset | Relevant | N/A | |||||
| CSQA | 70.8 | 73.4 | 71.2 | 72.9 | 72.6 | 74.6 | 74.1 |
| MBPP | 58.2 | 59.8 | 60.6 | 59.6 | 60.2 | 62.0 | 61.4 |
| GPQA | 31.6 | 34.4 | 33.7 | 33.1 | 32.6 | 35.8 | 36.2 |
@@
Our experiments and analysis so far used GPT-3.5 as the LLM, which is closed-source and gets updated over time. To verify whether the observations and conclusions are consistent across different models and additionally for reproducibility, we extend the experiments to Llama-2-Chat (Touvron et al., 2023). Specifically, we use vLLM to serve a Llama-2-70b-Chat model for experiments and report the results of different methods/variants on GSM8K in Table 9 and Table 10. We can draw similar observations: (i) self-generated relevant examples underperform all types of irrelevant ones, (ii) ‘GPT4-Calibration’ consistently outperforms the other two variants, and (iii) and perform better than other approaches, demonstrating that the conclusions can be generalized to different models.
We further conduct experiments with Llama-3-8B-Instruct, Llama-3.1-8B-Instruct (Dubey et al., 2024) and Qwen2.5-14B-Instruct (Yang et al., 2024). The results reported in Table 9 demonstrate the generalizability of the conclusions across different model types and scales. In addition, since investigating analogical reasoning requires LLMs to self-generate different types of problems, we only experiment with instruction-tuned LLMs to ensure that they can follow the given instructions.
Generalization to Different Tasks
To test the generalizability of our findings beyond the math domain, we further conduct experiments on CommonsenseQA (commonsense reasoning) (Talmor et al., 2019), MBPP (code generation) (Austin et al., 2021) and GPQA (question answering of very hard questions) (Rein et al., 2024). The comparison between different methods is shown in Table 11, which demonstrates that our findings can be generalized to different types of tasks.
Comparison Beyond Analogical Reasoning
We consider two widely used methods Self-consistency (Wang et al., 2023b) and Auto-CoT (Zhang et al., 2023b), and compare our designed approaches with them on GSM8K using Llama-3.1-8B-Instruct. For Self-consistency, we employ 5 decoding paths for majority voting. The results reported in Table 12 demonstrate that our methods can also outperform other baselines beyond analogical reasoning.
In addition, we show the robustness to prompt format, the effect of the number of demonstrations, more analysis on and , the results of repeating problems and explicitly controlling the semantics of generated examples in Appendix A.7 A.11, respectively.
| Relevant | Self-consistency | Auto-CoT | ||
| 74.8 | 77.6 | 75.9 | 80.2 | 81.0 |
@@ 5 Conclusion
This work has several limitations. First, due to the inference cost of ChatGPT, we conduct experiments on subsets of the test data for mathematical reasoning tasks. Besides, we include 6 datasets requiring different reasoning capabilities in this work. A further improvement could be to explore more diverse types of tasks.
@@ A.2 Detailed Results for Different Random Seeds
| Average | Relevant | ||||||
| N/A | |||||||
| Method | Relevant | N/A | |||||
| GPT-4o-mini | 90.7 | 91.9 | 92.6 | 92.3 | 93.2 | 94.2 | 94.5 |
A.3 Results with GPT-4o-mini
We conduct experiments with GPT-4o-mini on GSM8K and present the results in Table 15, verifying the generalizability of our findings to GPT-4o-mini.
| Relevant | N/A | Oracle | ||||
| GSM8K | 0.50 | 0.16 | 0.28 | 0.19 | 0.08 | 0.62 |
A.4 Decomposition of Relevance
The relevance can be further separated into semantic relevance and procedure (reasoning steps) relevance. Our analysis in Q1-1 has demonstrated that semantic relevance does not matter. To investigate the importance of procedure relevance, we perform a similar analysis. Specifically, we compute the average cosine similarity between the rationales of the generated examples and the rationale of the query to quantitatively measure their relevance. The results on GSM8K are reported in Table 16, which highlight that procedure relevance is not the key factor for analogical reasoning performance on mathematical reasoning tasks.
A.5 Prompts and Outputs for Example Generation
A.6 Guided Problem Generation
@@
| Relevant | N/A | ||||
| 71.2 | 74.9 | 75.3 | 75.9 | 74.3 | |
| 72.0 | 75.2 | 74.7 | 76.2 | 75.5 |
| Number | Relevant | N/A | |||
| 3 | 73.1 | 77.3 | 75.0 | 75.3 | 75.5 |
| 5 | 71.5 | 75.5 | 75.1 | 76.3 | 75.3 |
| 56.3 | 50.9 | 56.4 | 51.3 |
| Method | Task | ||
| GSM8K | MATH | Average | |
| 75.7 | 36.8 | 56.3 | |
| 73.8 | 36.2 | 55.0 | |
A.7 Robustness to Prompt Format
To verify the robustness of different methods to prompt format, we experiment with two new prompts paraphrased from the original one by GPT-4 and present the results on GSM8K in Table 17. We also observe better performance with irrelevant examples than relevant ones, showing the robustness.
A.8 Different Numbers of Demonstrations
A.9 More Analysis on and
Our designed method generates correct and relevant examples, and generates correct and irrelevant examples. From the results in Table 8, we can see that and achieve similar average performance, demonstrating that relevance does not matter.
We further change the correct answers of the demonstrations in and to random answers, obtaining and . Obviously, generates incorrect and relevant examples, and generates incorrect and irrelevant examples. The comparison between these four methods in Table 19 further supports our claim that the key factor influencing the performance on mathematical reasoning is the accuracy of self-generated examples rather than their relevance.
A.10 Repeating Problems
| Relevant | N/A | Similar and Correct | Different and Correct | |
| 74.8 | 77.3 | 78.4 | 80.3 | 80.6 |
A.11 Explicit Semantic Control
We explore explicitly controlling the semantics of generated examples (including both problems and reasoning paths) on GSM8K using Llama-3.1-8B-Instruct. Specifically, we investigate the following two approaches: (i) prompting the model to generate semantically similar and correct examples, and (ii) prompting the model to generate semantically different and correct examples. The results reported in Table 21 further verify the correctness of our conclusions.